OpenStudy (anonymous):

___sinb=1/2[sin(a+b) - sin(a-b)]

jimthompson5910 (jim_thompson5910):

hint: page 2 http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (anonymous):

\[\sin \left( a+b \right)=\sin a \cos b+\cos a \sin b\] \[\sin \left( a-b \right)=\sin a \cos b-\cos a \sin b\] subtract and get the answer